# Arithmetic aspects of discrete periodic Toda flows

### v2

@@ Abstract
We construct a new algebraic linearization of the discrete periodic Toda flow by using Mumford’s algebraic description of the Jacobian of a hyperelliptic curve. In particular, the discrete periodic Toda flow can be expressed in terms of the famous Gauß composition law for quadratic forms adapted to the framework of hyperelliptic curves by Cantor. One surprising consequence of our approach is a new integrality property for the discrete periodic Toda flow which leads to a $p$-adic description of the closely related periodic box-ball flow, which has very surprising connections to number theory.
1 Introduction
1.1 Motivation
The periodic box-ball system is a cellular automaton (or a tropical integrable system) first introduced in [19] by Tokihiro and Yura. It consists of $N>2$ boxes on a circle, each box being either empty (00) or containing a ball ($1$), with the total number of balls smaller than $N/2$, i.e., the state space is given by $\mathcal{V}_{N}=\{y\in\{0,1\}^{N}|\sum_{i=1}^{N}y_{i}<N/2\}$. The periodic box-ball flow $\mathcal{B}:\mathcal{V}_{N}\to\mathcal{V}_{N}$ moves every ball to the nearest empty box to the right, where boxes containing a ball are forbidden. Let us explain the flow using the example $y_{13}=1100011000101\in\mathcal{V}_{13}$:
[…]
As the state space $\mathcal{V}_{N}$ is finite and the flow $\mathcal{B}$ invertible, for every state $y\in\mathcal{V}_{N}$ there is a smallest number called the period $\mathrm{per}(y)\geq 1$ such that $\mathcal{B}^{\mathrm{per}(y)}(y)=y$. In our example we have $\mathrm{per}(y_{13})=273$.
Let us now recall Chebyshev’s prime counting function
$\psi(n)=\sum_{k\geq 1}\sum_{p^{k}\leq n}\mathrm{log}(p)=\mathrm{log}(\mathrm{lcm}(1,2,\dots,n)).$ (1.2)
Very surprisingly, using combinatorial methods, Tokihiro and Mada showed in [14] that (roughly speaking) for all $N>0$ there exists a state $y_{N}\in\mathcal{V}_{N^{2}}$ such that
$\mathrm{per}(y_{N})\approx\mathrm{exp}(\psi(N)).$ (1.3)
In particular, as explained in [14], this implies the fascinating fact that the dynamics of periodic box-ball systems encode the Riemann hypothesis, which can be formulated in terms of certain asymptotic behaviour of Chebyshev’s prime counting function $\psi$.
@@ 1.1 Motivation
This motivates the study of the geometric and arithmetic structures underlying the periodic box-ball system. For this we study the famous discrete periodic Toda system, a closely related discrete integrable system, see [11, 8]. To connect the two, we recall that the periodic box-ball system is essentially equivalent to the tropical periodic Toda system which can in turn be lifted to the discrete periodic Toda system. The slogan to keep in mind is that the periodic box-ball flow can be recovered from the discrete periodic Toda flow.
1.2 Overview of our results
@@
The main idea of the present paper is to construct a new linearization of the discrete periodic Toda flow using Mumford’s algebraic description of the Jacobian $\mathrm{Jac(\mathcal{C})}$ of the corresponding spectral curve $\mathcal{C}$. The addition law on $\mathrm{Jac(\mathcal{C})}$ is given by Gauß famous composition law for quadratic forms adapted to the setting of hyperelliptic curves by Cantor [1]. In particular, our linearization describes the discrete periodic Toda flow in terms of the Gauß composition law for hyperelliptic curves, a beautiful and surprising fact. On $\mathrm{Jac(\mathcal{C})}$ the flow translates to addition with the element $\mathfrak{T}\in\mathrm{Jac(\mathcal{C})}$, as explained in Theorem 2.2, one of our main results.
The main ingredient used for the linearization is given by the famous eigenvector map $\Psi_{\mathcal{C}}$ (see Definition 2.7).
The fact that the discrete periodic Toda system can be linearized on the Jacobian of the associated spectral curve is well-known, in particular the theory of $\theta$-functions on the analytic Jacobian has been used to linearize the flow (see [11, 9]).
The advantages of our new algebraic approach are manifold. First of all, we obtain a new and very transparent description of the inverse Abel-Jacobi map (see Proposition 2.1). The central role of the fundamental solution of the Pell-Abel equation of $\mathcal{C}$, which corresponds to a torsion point $\mathfrak{pa}\in\mathrm{Jac(\mathcal{C})}$, becomes apparent. Not only does it correspond to the cyclic shift map (see Theorem 2.1), it can also be used to describe the discrete periodic Toda flow (see Proposition 2.2).
Traditionally the discrete periodic Toda flow has be studied over $\mathbb{R}$ or $\mathbb{C}$, but our treatment makes it clear, that the discrete periodic Toda flow over the rational function field $\mathbb{Q}(T)$ is more fundamental. For example, it nicely recovers the tropical periodic Toda and periodic box-ball flows (see Proposition 3.2).
Even more interestingly we uncover a new integrality property (see Theorem 4.1) which shows that the discrete periodic Toda flow is defined over local rings, such as $\mathbb{Z}[T]_{(T)}$ or $\mathbb{Z}_{p}$. In particular we obtain a $p$-adic periodic Toda flow which recovers the periodic box-ball flow (see Corollary 4.1). This is another main result of our paper.
This opens the way to use $p$-adic methods to study the discrete periodic Toda flow. For example, using $p$-adic differential equations (see [3]) it should be possible to extend Cantor’s theory of division polynomials for imaginary hyperelliptic curves (see Theorem 2.3) to the framework of real hyperelliptic curves, which appear as spectral curves of discrete periodic Toda systems. Surprisingly, this seems to be an open problem in the literature.
1.3 Summary
In section 2) we introduce the discrete periodic Toda flow and the different descriptions of the state space. We explain the notion of associated spectral curve, explain Mumford’s algebraic Jacobian and its addition law (in the context of real hyperelliptic curves) given by the famous Gauß composition law. The central ingredient is the eigenvector map $\Psi_{\mathcal{C}}:\mathcal{R}_{\mathcal{C}}\to\mathrm{Jac(\mathcal{C})}$ on which the linearization procedure is based. We explain fundamental properties of $\Psi_{\mathcal{C}}$, such as its inverse Abel-Jacobi map, the compatibility result with the cyclic shift map $\sigma$, certain involutions and most importantly the discrete periodic Toda flow. We also explain how the discrete periodic Toda flow is related to the shift map $\sigma$. In the end we explain Cantor’s theory of division polynomials (for imaginary hyperelliptic curves).
In section 3) we explain how one can recover the periodic box-ball and the tropical periodic Toda flows from the discrete periodic Toda flow. Here it is important to work over $\mathbb{Q}(T)$ in our set-up. We also indicate a tropical version of Mumford’s Jacobian.
In section 4) we explain our integrality results and end the paper with
section 5) where we give an outlook and ask some questions.
In the appendix we relate our approach to the monodromy matrix approach from [11].
1.4 Notations and conventions
We define the natural numbers by =$\mathbb{N}=\{0,1,2,\cdots\}$ and let $R$ be a nice ring in characteristic 0 over which our set-up works. We always assume that $R$ embeds into its field of fractions $K$. Traditionally, fields like $\mathbb{R}$ or $\mathbb{C}$ are used, but in our set-up the field of rational functions $\mathbb{Q}(T)$ will appear prominently. Later, local rings, such as $\mathbb{Z}[T]_{(T)}$ and $\mathbb{Z}_{(p)}\subset\mathbb{Z}_{p}$ will appear as well [^2] One can safely assume that $R$ is always one of the before mentioned rings..
Let $M\in R^{n\times n}$ be a quadratic matrix. We denote by $|M|$ its determinant.
The matrix
$M_{i,j}$ (1.4)
denotes the matrix obtained by deleting the $i$-th column and $j$-th row of $M$. We write
${}^{i}M_{j}\in R^{n-(i+j)\times n-(i+j)}$ (1.5)
for the matrix obtained by deleting the first $i$ rows and columns and by deleting the last $j$ rows and columns.
Usually, we will suppress a 00 from the notation, i.e., e.g., we write $M_{j}={}^{0}M_{j}$.
Finally, we denote by
$\mathcal{M}=M+xE_{n}$ (1.6)
the characteristic matrix of $M$.
For a polynomial $P(x_{1},..,x_{n})\in R[x_{1},..,x_{n}]$ we write
$P_{x_{j}}=P(x_{1},..,x_{j-1},0,x_{j+1},..,x_{n}).$ (1.7)
2 Linearization of the discrete periodic Toda flow
@@
2.1 The discrete periodic Toda flow
The discrete periodic Toda flow over $R$ (for $n>1$) is the rational map
$\mathfrak{T}:R^{2n}\dashrightarrow R^{2n},$ (2.1)
$(\mathbb{I}^{t},\mathbb{V}^{t})=(I_{1}^{t},\dots,I_{n}^{t},V_{1}^{t},\dots,V_{n}^{t})\longmapsto(\mathbb{I}^{t+1},\mathbb{V}^{t+1}),$ (2.2)
@@ 2.1 The discrete periodic Toda flow
$\begin{split}I^{t+1}_{i}&=I_{i}^{t}+V_{i}^{t}-V^{t+1}_{i-1},\\ V_{i}^{t+1}&=\frac{I^{t}_{i+1}V_{i}^{t}}{I_{i}^{t+1}},\end{split}$ (2.3)
@@
$0\neq\prod_{i=1}^{n}I^{t}_{i}\neq\prod_{i=1}^{n}V^{t}_{i}\neq 0$ (2.5)
Under the above assumption, using Proposition 2.1, one can show
$I^{t+1}_{i}=I^{t}_{i+1}\tfrac{I_{i}^{t}I_{i-1}^{t}\cdots I^{t}_{i+2}+V^{t}_{i}I^{t}_{i-1}\cdots I^{t}_{i+2}+\dots+V^{t}_{i}V^{t}_{i-1}\cdots V^{t}_{i+2}}{I_{i-1}^{t}I_{i-2}^{t}\cdots I_{i+1}^{t}+V^{t}_{i-1}I^{t}_{i-2}\cdots I^{t}_{i+1}+\dots+V^{t}_{i-1}V^{t}_{i-2}\cdots V^{t}_{i+1}}$ (2.6)
@@
$V^{t+1}_{i}=V_{i}^{t}\tfrac{I_{i-1}^{t}I_{i-2}^{t}\cdots I_{i+1}^{t}+V^{t}_{i-1}I^{t}_{i-2}\cdots I^{t}_{i+1}+\dots+V^{t}_{i-1}V^{t}_{i-2}\cdots V^{t}_{i+1}}{I_{i}^{t}I_{i-1}^{t}\cdots I^{t}_{i+2}+V^{t}_{i}I^{t}_{i-1}\cdots I^{t}_{i+2}+\dots+V^{t}_{i}V^{t}_{i-1}\cdots V^{t}_{i+2}}.$ (2.7)
2.2 Jacobi matrices

### v3

@@ Abstract.
We construct a new algebraic linearization of the discrete periodic Toda flow by using Mumford’s algebraic description of the Jacobian of a hyperelliptic curve. In particular, the discrete periodic Toda flow can be expressed in terms of the famous Gauß composition law for quadratic forms adapted to the framework of hyperelliptic curves by Cantor. One remarkable consequence of our approach is a new integrality property for the discrete periodic Toda flow. This leads, in particular, to a $p$-adic description of the closely related periodic box-ball flow, which has very surprising connections to number theory.
1. Introduction
1.1. Motivation
The periodic box–ball system is an integrable cellular automaton introduced by Tokihiro and Yura [19]. It evolves configurations of $N>2$ boxes arranged on a circle, where each box is either empty (0) or contains a ball (1), under the constraint that the total number of balls is <N/2<N/2. Thus the phase space is
$\mathcal{V}_{N}=\Bigl\{y\in\{0,1\}^{N}\ \Big|\ \sum_{i=1}^{N}y_{i}<\tfrac{N}{2}\Bigr\},$
and the time evolution $\mathcal{B}:\mathcal{V}_{N}\to\mathcal{V}_{N}$ moves each ball to the nearest empty box to the right, with occupied boxes forbidden. Since $\mathcal{V}_{N}$ is finite and $\mathcal{B}$ is invertible, every state $y\in\mathcal{V}_{N}$ is periodic; we write $\mathrm{per}(y)\geq 1$ for the minimal period.
Figure 1. Evolution of the state $y_{13}=1100011000101\in\mathcal{V}_{13}$ with $\mathrm{per}(y_{13})=273$.
A remarkable discovery of Tokihiro and Mada shows that the distribution of these periods is linked to deep number theory: roughly speaking, for each $N$ one can find a state whose period grows like $\exp(\psi(N))$, where $\psi(k)=\mathrm{log}(\mathrm{lcm}(1,2,\dots,k))$ is the Chebyshev’s famous second function. As explained in [14], this viewpoint allows one to formulate the Riemann hypothesis in terms of asymptotic properties of the periodic box–ball dynamics.
@@ 1.1. Motivation
These connections motivate a systematic study of the geometric and arithmetic structures underlying periodic box–ball systems. A key bridge is provided by the tropical periodic Toda system, which is closely related to the box–ball system and admits a lift to the discrete periodic Toda system. In this paper we therefore study the discrete periodic Toda flow and show how the periodic box–ball flow can be recovered from it.
1.2. Overview of results.
@@
The main idea of this paper is to construct a new algebraic linearization of the discrete periodic Toda flow using Mumford’s description of the Jacobian $\mathrm{Jac(\mathcal{C})}$ of the associated hyperelliptic spectral curve $\mathcal{C}$. The group law on $\mathrm{Jac(\mathcal{C})}$ is implemented by Cantor’s adaptation of Gauß’ composition law for quadratic forms to the hyperelliptic setting.
The core technical input is the eigenvector (or algebraic Abel-Jacobi) map
$\Psi_{C}:\mathcal{R}_{\mathcal{C}}\longrightarrow\mathrm{Jac(\mathcal{C})},$
which assigns to a Toda state on the isolevel set $\mathcal{R}_{\mathcal{C}}$ a Mumford representation in $\mathrm{Jac}(\mathcal{C})$. While it is classical that discrete periodic Toda flows linearize on Jacobians (typically via $\theta$-functions on the analytic Jacobian), our approach stays entirely algebraic and yields several new structural and arithmetic consequences:
  • Explicit inverse Abel–Jacobi map. We obtain a transparent algebraic description of $\Psi_{\mathcal{C}}^{-1}$ (Proposition 2.1), giving explicit recovery formulas for Toda variables from a Mumford representation.
  • Cyclic shift $\sigma$ and torsion. The cyclic shift map $\sigma$ on the Toda phase space corresponds under $\Psi_{\mathcal{C}}$ to translation by a distinguished torsion point $\mathfrak{pa}\in\mathrm{Jac}(\mathcal{C})$ arising from a fundamental solution of the Pell–Abel equation of $\mathcal{C}$ (Theorem 2.1).
  • Linearization of the discrete Toda flow. On $\mathrm{Jac}(C)$ the Toda time evolution is translation by an element $\mathfrak{T}\in\mathrm{Jac}(\mathcal{C})$ (Theorem 2.2), giving a concrete description of the discrete periodic Toda flow in terms of the Gauß–Cantor composition law.
@@
  • Discrete Toda flow via $\sigma$. The discrete periodic Toda flow can be expressed in terms of the cyclic shift $\sigma$ (Proposition 2.2).
  • Why $\mathbb{Q}(q)$ is natural. Although Toda flows are traditionally studied over $\mathbb{R}$ or $\mathbb{C}$, our framework highlights that working over the rational function field $\mathbb{Q}(q)$ is much more natural and fundamental. In particular, via the $q$-adic valuation this setting recovers both the tropical periodic Toda flow and the periodic box–ball flow (Proposition 3.2 & Theorem 3.1).
  • Integrality and $p$-adic Toda flow. A common belief in the field of integrable systems is that the discrete periodic Toda flow is only defined over a field.
    Surprisingly, we prove that the discrete periodic Toda flow is defined over local rings such as $\mathbb{Z}[q]_{(q)}$ (Theorem 4.1). Specializing $q\mapsto p$ then yields $p$-adic Toda flows and recovers the periodic box–ball flow via the $p$-adic valuation (Corollary 4.1).
Finally, this arithmetic viewpoint suggests that $p$-adic methods (for instance those used in [3]) may help extend Cantor’s division-polynomial theory from imaginary hyperelliptic curves to the real hyperelliptic curves that naturally occur as Toda spectral curves (cf., [18]).
1.3. Organization of the paper.
Section 2 introduces the discrete periodic Toda flow, the associated spectral curve, Mumford’s algebraic Jacobian and the Gauß–Cantor composition law, and develops the eigenvector map $\Psi_{\mathcal{C}}$ together with its inverse and its compatibilities (shift, involutions, and time evolution).
Section 3 explains how periodic box–ball and tropical periodic Toda dynamics are recovered from the discrete periodic Toda flow, emphasizing the role of $\mathbb{Q}(q)$ and valuations.
Section 4 proves the integrality result and derives $p$-adic consequences, and Section 5 collects questions and outlook. Appendix A relates our approach to the monodromy-matrix viewpoint. Appendix B reviews Cantor’s division polynomials (in the imaginary case) and explains how they can be used to compute multiples of the translation element on $\mathrm{Jac}(\mathcal{C})$ efficiently. Appendix C discusses a tropical analogue of Mumford’s Jacobian, including a tropicalized inverse Abel–Jacobi map and the induced tropical linearization picture.
As our paper relates two different areas of research (integrable systems and number theory) we tried to make the paper accessible to researchers from both fields. Further, we have checked all computations with the help of SageMath.
1.4. Acknowledgements
We wish to express our sincere gratitude to the (anonymous) referees for carefully reading the manuscript and for their valuable comments, which led to considerable improvements.
1.5. Notation and conventions.
Our paper involves several layers of objects (Toda states, Jacobi matrices, spectral curves, Jacobians, and their tropical counterparts). To keep symbols unambiguous we fix our conventions once and for all and collect the most frequently used notation in the following table.
Symbol Meaning First
General conventions
$\mathbb{N}$ $\{0,1,2,\dots\}$
$n$ size of Toda system §2.1
NN number of boxes in the periodic box–ball system §1.1
$R\subset K$ base ring [^1] One can safely assume that $R$ is always one of the aforementioned rings.(char. 0) with fraction field $K$; key examples include fields: $\mathbb{R},\mathbb{Q}(q)$ and local rings: $\mathbb{Z}[q]_{(q)},\mathbb{Z}_{p}$ §2.1
Linear algebra
$M\in R^{n\times n}$ matrix with entries in $R$ §2.2
$\mathcal{M}=M+xE_{n}$ characteristic matrix of $M$ (with $E_{n}$ the unit matrix) §2.3
$|M|\in R$, $|\mathcal{M}|\in R[x]$ determinant of $M$ and $\mathcal{M}$ §2.3
$M_{i,j}\in R^{(n-1)\times(n-1)}$ minor: delete the $i$-th row and $j$-th column §2.3
$^{i}M_{j}\in R^{(n-i-j)\times(n-i-j)}$ submatrix: delete first $i$ rows/cols and last $j$ rows/cols §2.3
$M_{j},\ {}^{i}\!M$ abbreviations $M_{j}={}^{0}\!M_{j}$ and $^{i}\!M={}^{i}\!M_{0}$ Lem. 2.1
$p_{x}$ evaluation at $x=0$: $p_{x}=p(0)$ for $p\in R[x]$ Rem. 2.5
$f|_{I_{j}=0}$, $f|_{V_{j}=0}$ set $I_{j}=0$ in $f\in R[I_{1},\dots,I_{n},V_{1},\dots,V_{n}][x]$ (similarly $f|_{V_{j}=0}$) Lem. 2.4
$f_{I_{j}},\ f_{V_{j}}$ shorthand for $f|_{I_{j}=0}$ and $f|_{V_{j}=0}$ Lem. 2.4
Discrete periodic Toda
$(\mathbb{I}^{t},\mathbb{V}^{t})\in R^{2n}$ Toda state at time $t$, with components $I_{i}^{t},V_{i}^{t}$; indices are periodic $i=i+n$ §2.1
$\mathcal{T}:R^{2n}\dashrightarrow R^{2n}$ Toda time flow/step: $(\mathbb{I}^{t+1},\mathbb{V}^{t+1})=\mathcal{T}(\mathbb{I}^{t},\mathbb{V}^{t})$ (2.1)
$\sigma:R^{2n}\to R^{2n}$ cyclic shift on indices $i\mapsto i+1$ (with $i=i+n$) and induced action: $(\mathbb{I}^{t,\sigma},\mathbb{V}^{t,\sigma})=\sigma(\mathbb{I}^{t},\mathbb{V}^{t})$ Def. 2.2
$^{\dagger}:R^{2n}\to R^{2n}$ dagger involution and induced action: $(\mathbb{I}^{t,\dagger},\mathbb{V}^{t,\dagger})=(\mathbb{I}^{t},\mathbb{V}^{t})^{\dagger}$ Def. 2.3
$z$ invertible spectral parameter in $L_{t}$ §2.2
$L_{t}$, $\overline{L}_{t}$, $L^{\sigma}_{t}$, $L^{\dagger}_{t}\in R(z)^{n\times n}$ periodic Jacobi/Lax matrices associated to Toda states $(\mathbb{I}^{t},\mathbb{V}^{t})$, $(\mathbb{I}^{t+1},\mathbb{V}^{t+1})$, $(\mathbb{I}^{t,\sigma},\mathbb{V}^{t,\sigma})$, $(\mathbb{I}^{t,\dagger},\mathbb{V}^{t,\dagger})$ §2.2
$\mathcal{L}_{t}$, $\overline{\mathcal{L}}_{t}$, $\mathcal{L}_{t}^{\sigma}$, $\mathcal{L}_{t}^{\dagger}\in R(z)[x]^{n\times n}$ characteristic matrices of $L_{t}$, $\overline{L}_{t}$, $L^{\sigma}_{t}$, $L_{t}^{\dagger}$ §2.3
$\sigma_{i,j}$, $s_{i}$, $t_{i}$ involutions $R^{2n}\to R^{2n}$ on Toda states Def. 2.5
Spectral curve and invariants
$\mathcal{C}=z|\mathcal{L}_{t}|$ spectral curve $z^{2}+h(x)z+f=0$ of genus g=n1g=n-1 (Toda flow invariant) §2.3
$h(x)\in R[x],\ f\in R$ coefficients defining $\mathcal{C}$ (Toda flow invariants) §2.3
$I_{\Pi},\ V_{\Pi}$ Toda flow invariants $I_{\Pi}=\prod_{i=1}^{n}I_{i}^{t}$, $V_{\Pi}=\prod_{i=1}^{n}V_{i}^{t}$ §2.3
$u^{t}$, $v^{t}$, $w^{t}\in R[x]$ polynomials extracted from $\mathcal{L}_{t}$ Def. 2.1
$\overline{u}^{t}$, $\overline{v}^{t}$, $\overline{w}^{t}\in R[x]$ polynomials extracted from $\overline{\mathcal{L}}_{t}=\mathcal{L}_{t+1}$ Def. 2.1
$u^{t,\sigma}$, $v^{t,\sigma}$, $w^{t,\sigma}\in R[x]$ polynomials extracted from $\mathcal{L}_{t}^{\sigma}$ Def. 2.4
$u^{t,\dagger}$, $v^{t,\dagger}$, $w^{t,\dagger}\in R[x]$ polynomials extracted from $\mathcal{L}_{t}^{\dagger}$ Def. 2.4
Linearization on Mumford Jacobian
$\mathrm{Jac(\mathcal{C})}$ Jacobian of $\mathcal{C}$; elements written in balanced Mumford form $([P,Q],d)$ §2.6
$[P,Q]$ Mumford representation (with $P\mid(Q^{2}+hQ+f)$ and $\deg P\leq g$) §2.6.1
$\boxplus$ addition on $\mathrm{Jac(\mathcal{C})}$ via Gauß–Cantor composition law §2.7
$\mathcal{R}_{\mathcal{C}}\subset R^{2n}$ isolevel set of Toda states with fixed spectral curve $\mathcal{C}$ Def. 2.6
$\Psi_{\mathcal{C}}:\mathcal{R}_{\mathcal{C}}\to\mathrm{Jac}(\mathcal{C})$ eigenvector (or algebraic Abel-Jacobi) map $(\mathbb{I}^{t},\mathbb{V}^{t})\mapsto([u^{t},v^{t}],0)$ Def. 2.7
$\mathfrak{T}\in\mathrm{Jac}(C)$ translation element inducing Toda time flow on $\mathrm{Jac}(C)$ Thm. 2.2
$\mathfrak{pa}\in\mathrm{Jac(\mathcal{C})}$ Pell–Abel torsion point $\infty_{+}-\infty_{-}]\in\mathrm{Jac}(\mathcal{C})$, used to describe action of $\sigma$ and on $\mathrm{Jac(\mathcal{C})}$ §2.6.3
Tropical periodic Toda / periodic box–ball
$\mathcal{B}:\mathcal{V}_{N}\to\mathcal{V}_{N}$ periodic box–ball flow on $\mathcal{V}_{N}\subset\{0,1\}^{N}$. §1.1
$\mathcal{T}^{\mathrm{trop}}:\mathbb{R}^{2n}\dashrightarrow\mathbb{R}^{2n}$ tropical periodic Toda flow on $\mathbb{R}^{2n}$ (often restricted to $\mathbb{N}^{2n}$) §3.2
$\eta:\mathcal{V}_{N}\longrightarrow\bigsqcup_{n=1}^{\lfloor N/2\rfloor}\mathbb{N}^{2n}$ embedding from box–ball states into tropical Toda states. §3.3
$\mathfrak{I}_{q}:\mathbb{N}^{2n}\to\mathbb{Q}(q)^{2n}$ (algebraic) lift from tropical to discrete periodic Toda (3.2)
$\mathcal{T}_{p}:\mathbb{Z}_{p}^{2n}\dashrightarrow\mathbb{Z}_{p}^{2n}$ $p$-adic periodic Toda flow on $\mathbb{Z}_{p}^{2n}$ §3.2
Table 1. List of most frequently used symbols and conventions.
From Section 2 onward we often suppress the time superscript $t$ on $(\mathbb{I}^{t},\mathbb{V}^{t})$ and on the associated objects, e.g., $L_{t},u^{t},v^{t},w^{t}$, when no confusion can arise. Whenever a statement compares different times (typically $t$ and $t+1$), we write the time superscripts explicitly or equivalently use $u=u^{t}$ and $\overline{u}=u^{t+1}$.
2. Linearization of the discrete periodic Toda flow
2.1. The discrete periodic Toda flow
The discrete periodic Toda flow over $R$ (for $n>1$) is the rational map on Toda states
$\mathcal{T}:R^{2n}\dashrightarrow R^{2n},$
$(\mathbb{I}^{t},\mathbb{V}^{t})=(I_{1}^{t},\dots,I_{n}^{t},V_{1}^{t},\dots,V_{n}^{t})\longmapsto(\mathbb{I}^{t+1},\mathbb{V}^{t+1}),$
@@ 2.1. The discrete periodic Toda flow
(2.1) $\begin{split}I^{t+1}_{i}&=I_{i}^{t}+V_{i}^{t}-V^{t+1}_{i-1},\\ V_{i}^{t+1}&=\frac{I^{t}_{i+1}V_{i}^{t}}{I_{i}^{t+1}},\end{split}$
@@
$0\neq\prod_{i=1}^{n}I^{t}_{i}\neq\prod_{i=1}^{n}V^{t}_{i}\neq 0.$
Under the above assumption, one can show (cf., [11])
$I^{t+1}_{i}=I^{t}_{i+1}\tfrac{I_{i}^{t}I_{i-1}^{t}\cdots I^{t}_{i+2}+V^{t}_{i}I^{t}_{i-1}\cdots I^{t}_{i+2}+\dots+V^{t}_{i}V^{t}_{i-1}\cdots V^{t}_{i+2}}{I_{i-1}^{t}I_{i-2}^{t}\cdots I_{i+1}^{t}+V^{t}_{i-1}I^{t}_{i-2}\cdots I^{t}_{i+1}+\dots+V^{t}_{i-1}V^{t}_{i-2}\cdots V^{t}_{i+1}}$
@@
$V^{t+1}_{i}=V_{i}^{t}\tfrac{I_{i-1}^{t}I_{i-2}^{t}\cdots I_{i+1}^{t}+V^{t}_{i-1}I^{t}_{i-2}\cdots I^{t}_{i+1}+\dots+V^{t}_{i-1}V^{t}_{i-2}\cdots V^{t}_{i+1}}{I_{i}^{t}I_{i-1}^{t}\cdots I^{t}_{i+2}+V^{t}_{i}I^{t}_{i-1}\cdots I^{t}_{i+2}+\dots+V^{t}_{i}V^{t}_{i-1}\cdots V^{t}_{i+2}}.$
Remark 2.1.
This also follows easily from our new inverse algebraic Abel-Jacobi map, using Proposition 2.1.
2.2. Jacobi matrices